OpenStudy (anonymous):

Use trigonometric identities to transform the left side of the equation into the right side (0 < θ < π/2). tan θ cosθ= sinθ

hartnn (hartnn):

on left side, use the standard relation that \(\Large \tan \theta = \dfrac{\sin \theta }{\cos \theta }\)

OpenStudy (anonymous):

This one's for tan theta cos theta = sin theta?

hartnn (hartnn):

yes, replace tan theta on left side by sin theta/ cos theta notice that cos theta gets cancelled what remains is sin theta which is your right side. thats it!

OpenStudy (anonymous):

Oh, thank you! Sorry, I don't really know how to use this thing yet. By the way, I dig your profile pic!:)

hartnn (hartnn):

you're welcome ^_^ did you like it ? :)

OpenStudy (anonymous):

Yes!:3

hartnn (hartnn):

thanks you :)

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

Could you help me with another problem?

OpenStudy (anonymous):

A steel plate has the form of one-fourth of a circle with a radius of 60 centimeters. Two two-centimeter holes are to be drilled in the plate positioned as shown in the figure. Find the coordinates of the center of each hole.

OpenStudy (anonymous):

 Sorry for my bad artistic skills. I hope you still understand it.